OpenStudy (anonymous):

Find the width of a rectangle with a perimeter of 90 and a length of 15. a.90 b.15 c.45 d.30

OpenStudy (anonymous):

The perimeter of a rectangle is the sum of all the sides of that rectangle..

OpenStudy (mathstudent55):



OpenStudy (anonymous):



OpenStudy (anonymous):

In a rectangle, AD = BC And AB = CD

OpenStudy (anonymous):

Now, Perimeter of the rectangle is given as: Perimeter = AB + BC + CD + DA Perimeter = AB + BC + AB + BC Perimeter = 2AB + 2BC Perimeter = 2(Length + Width)

OpenStudy (triciaal):

from the diagrams above you see the perimeter, P = 2( length + width) given P = 90 and length = 15 plug the numbers and solve for the width

OpenStudy (anonymous):

So, you are given Perimeter as 90 and length as 15.. So, \(90 = 2(15 + Width)\) \(45 = 15 + Width\) Subtract 15 from both the sides, and you will get your Width..

OpenStudy (anonymous):

thank you @waterineyes @triciaal

OpenStudy (anonymous):

You are welcome dear... :)

OpenStudy (triciaal):

yes I agree with D. you are welcome